This question will involve consideration of conditions observed on two different days. On Day One, the surface temperature is 25°C and the surface dewpoint is 15°C. On Day Two, the surface temperature is 25°C and the surface dewpoint temperature is 20°C. If cumulus clouds form on both days, on which of the two days is the cloudbase (i.e. the height of the bottom of the cloud) lower? Clearly explain the reasoning you used to arrive at your answer.
Answer #1

you can find out the cloud base by using the equation given below.

Cloud base = {(Surface Temperature - Dewpoint Temperature) /2.5 } x 1000

In words- First find the difference between surface temperature and Dewpoint Temperature (this is called 'spread', divide it with 2.5. The answer so obtained has to be multiplied by 1000. You will get the answer.

Things to be noted

> The height so obtained is in Feet

> You should add the elevation of the area with the answer to get the cloud base from sea level.

I will just workout the cloud base of Day 1

> spread = 25-15= 10; 10/2.5= 4; 4 x 1000 = 4000ft